# Let’s Influence Algorithms Together: How Millions of Fans Build Collective Understanding of Algorithms and Organize Coordinated Algorithmic Actions

### v1

@@ 4.2. Fan-Driven Folk Theories: The Foundation for Guiding Collective Algorithmic Actions (RQ2)
4.2.1. Fan-Driven Methods in Understanding Algorithms.
Before launching collective actions targeting algorithms, people must have a method to thoroughly understand the algorithms, allowing them to perceive whether their actions can effectively influence them and adjust their strategies accordingly based on this assessment. Core fans use the term ”decode” to describe the initial step in developing their understanding of algorithms. They see the algorithm as a black box; users are unaware of its exact operational logic and can only decode it through continuous trial and error. Eventually, this user-driven process creates a set of tutorials that can genuinely influence the algorithm’s output in fans’ perception.
@@
To develop a clear understanding of the algorithm, initially, core fans often gather basic tutorials from other fan groups, as these groups also engage in decoding algorithms and instructing their fans. This is because all fan groups are actively decoding various social media algorithms to increase the visibility of their idols. To engage more general fans, these algorithm manipulation tutorials are always made public. As previously mentioned, there is already a consensus among core and general fans that various interactions on social media significantly impact their idols. Then, it is important to understand which specific interactions can most effectively leverage recommendation algorithms to maximize the visibility of their idols.
@@
By observing patterns in algorithm behavior from various idol fanbases, core fans can formulate an initial strategy for their own group. For example, CF26 and her core fan colleagues examine tutorials from different idol fanbases, selecting and integrating them based on their experience. The core fan then guides general fans in improving the idol’s visibility through continual experimentation, eventually tailoring an algorithmic tutorial that best fits their fanbase. This is because fan groups are heterogeneous. The behavior and characteristics of fans from different idol groups can vary significantly, so core fans must carefully understand and adapt to the style of their corresponding general fan group. Moreover, since these are folk understandings, the algorithmic tutorials created by other fan groups are not necessarily effective. Core fans are often skeptical of tutorials from other groups and prefer to develop their own understanding through repeated experimentation and testing.
@@
Improvements in data performance are often perceptible only to core fans, rather than being quantifiable through objective social media metrics. They observe whether the manipulations lead to higher comment rankings and visibility than other fanbases. Recently, CF26 got information from a fan working at Sina (Weibo company) about how to make a topic become a trending topic. When a topic related to an idol enters the trending section by the recommended algorithm, it gains significant visibility, making this achievement extremely important for fans.
@@
Within a single topic, there are often many posts, but how to intervene so that the algorithm recommends the topic and pushes it into Weibo’s trending list, thereby gaining greater visibility, is a common concern for fans. CF26 mentioned that the new information from the Sina staff asks the fans to engage with the post displayed on the top of the topic, called “source of trend.” This means that interacting with the top posts within a topic—by giving them more retweets, likes, and comments—will make the topic more likely to enter the trending section rather than interacting with other, less prominent posts. When CF26 and her followers tried to make a new trend topic for their idol, they utilized this information to make a new tutorial, and engaged with the “source of trending.” After the topic successfully becomes a new trending topic, CF26 uploads a new post: “The new tutorial is so useful! It’s been trending for about two hours!” CF26 told us that she uses this period as a verification period to ensure the new tutorial makes sense.
@@
4.2.2. Expand Folk Understanding of Algorithms across Platforms in Different Nations.
If such collective action involves multiple countries, people must be able to understand the differences in platform and algorithm policies across these countries. It is important to note that experiences manipulating algorithms can vary across different nations. For instance, in the context of the Korean entertainment industry and social media, Korean fans often engage in a specific strategy of “building within a building,” meaning “commenting within comments,” a layered commenting approach to manipulating data performance.
CF26 highlighted the unique dynamics between Korean and Chinese social media, noting, ”The context in Korean social media differs from that in China, where the regulations are more stringent.” The ”die-hard fan” status, a Weibo feature in China, indicates a high level of interaction between followers and specific bloggers, requiring engagement for at least five days within the past 30 days. This status must be earned separately for each blogger. In China, the ”building within a building” commenting strategy is significantly influenced by the user’s status as a ”die-hard fan” and their VIP identity on Weibo. CF26 explained that if users meet both criteria, their fan group’s comments gain enhanced visibility. However, if they lack the ”die-hard fan” status, the strategy can actually have a detrimental effect.
Additionally, CF26 elaborated that most fans in China do not meet the ”die-hard fan” and VIP criteria. ”die-hard fan” represents a high frequency of interaction with a specific blogger, while achieving VIP status involves monetary contributions. General fans also find it difficult to consistently achieve ”die-hard fan” status across multiple bloggers. As a result, the ”building within a building” strategy may decrease visibility if those who cannot fulfill both requirements conduct algorithmic actions.
In contrast, Korean social media platforms usually do not have the same requirements for enhanced visibility, such as the ”die-hard fan” status and VIP identity. This means that anyone can use the ”building within a building” strategy in Korea to increase visibility without needing a high level of interaction or monetary contributions. The effectiveness of this strategy in Korea, regardless of a user’s status, highlights the different ways platform management policies are implemented in the two countries. While Chinese platforms have stricter rules that limit who can successfully use this strategy, Korean platforms allow all users to benefit from it equally.
CF43 agreed that in China, core fans typically advise against general fans using the ‘building within a building’ comment strategy. CF43 elaborated,
“This is because when general fans, who don’t meet the aforementioned criteria, attempt such commenting, the platform may classify it as an action of a ‘water army,’ perceived as disruptive to social media order. Once labeled as a ‘water army,’ the fan comment is de-prioritized and relegated to the bottom of the comment section, losing all visibility and rendering previous efforts futile.” (CF43)
This difference between the two countries also demonstrates that if core fans do not develop a precise understanding of the algorithms specific to each platform, their collective actions are unlikely to succeed. Without this knowledge, they cannot determine which behaviors will be effective or ineffective and may even risk causing negative consequences.
4.2.3. Expand Folk Understanding of Algorithm across Different Fan Bases.
@@
If this collective action involves different groups of people, it is essential to understand the unique characteristics of these groups and leverage these traits to effectively organize them. Consequently, the interpretation of algorithms and the formulation of ttutorials for algorithm manipulation by core fans vary significantly between cultures and fanbases, and are largely based on empirical observations.
@@
CF11, who manages a data group in an idol’s fan club, highlighted how fans from different countries, even those of the same idol, employ distinct strategies on platforms like YouTube. CF11 noted that view counts of fancams on YouTube are crucial for K-pop idols, and K-pop fans worldwide, regardless of their country, strive to increase these counts. However, CF11 pointed out a stark contrast in techniques: while international fan organizations advise fans to remain logged in while viewing, Chinese fans do not recommend this approach. CF11 further explained their understanding of YouTube’s view count algorithm: “YouTube’s algorithm limits the number of view counts to 3-5 plays per video per day from a single user. Users logged in can’t contribute more than five views. Therefore, to exceed this limit, one should not log in.” CF11 believes these different approaches stem from the varying user habits of the two fan bases,
“Chinese fans are used to repeatedly engaging in data labor, often exceeding five views per day, hence our preference for not logging in. In contrast, the international fan base, predominantly from Europe and the U.S., tends to view 3-5 times daily. For them, ensuring each view is counted becomes vital, making logged-in viewing more appropriate.” (CF11)
@@
4.2.4. Expand Folk Understanding of Algorithm across Different Levels of Algorithm Transparency.
@@
Since this is a collective action targeting algorithms, it must take into account the differences between the algorithms of various platforms, particularly in terms of transparency. Effectively leveraging platform transparency can enhance the efficiency of the action. Even when dealing with platforms with low transparency, people must strive to understand the underlying algorithmic mechanisms. While deciphering algorithms can be intricate and challenging, as seen with platforms like Weibo, some platforms offer more transparency. CF39 mentioned that many platforms provide official explanations of their recommendation systems, simplifying the process for core fans to create tutorials.
@@
For example, BiliBili, a popular platform in China, officially states that the same video can be counted 3-5 times per day per person for play data. Although full playback isn’t necessary for counting views, the completion rate in playing is a significant factor in the recommendation algorithm. This clarity gives fans straightforward tutorials for contributing to BiliBili’s play count.
@@
However, according to our participants, the criteria for BiliBili’s trending list are less clear for fans. Within the Chinese fandom, it’s rumored that a video must achieve 100k plays within 24 hours, along with a specific number of likes, shares, comments, and other metrics, to feature on the trending list algorithmically. Yet, there are instances where videos have made it to the trending list with fewer than 100k plays or exceeded this threshold without being featured. CF39 expressed frustration over this ambiguity, stating,
”We understand that activities like plays, likes, favorites, comments, and shares positively impact trending potential, but the exact mechanism is unclear. We cannot definitively influence the recommendation algorithm. Our only recourse is for fans to engage in data labor blindly and keep experimenting.” (CF39)
@@
We have described how core fans gradually develop a folk understanding of how algorithms operate on different platforms. It is important to note that while these understandings are not entirely conveyed to general fans, they serve as a crucial foundation for core fans to develop algorithm manipulation tutorials. All the core fans we interviewed acknowledged that without a clear understanding of platform algorithms, collective algorithmic action would be impossible. CF34, a core fan with over ten years of experience, explained to us,
”Ten years ago, we lacked understanding of algorithms and the idea of collective algorithmic action. Fans didn’t unite. Even after recognizing the importance of recommendation algorithms, it took time to learn how to coordinate effectively. Without clear direction, mobilizing millions of fans would have been impossible. Now, our actions are organized and disciplined, and fans know precisely what is expected of them, resulting in a significant impact.” (CF34)
@@
4.3. “Enable the Foolish to Learn, Motivate the Lazy to Act”: The Principles of Mobilizing Millions of Fans (RQ3)
@@
CF37 challenged the previous Chinese media’s portrayal of fans as a highly efficient, organized group that immediately responds to requests for collective algorithmic action. According to CF37, this depiction is misleading. CF37 explained that each act of data engagement—likes, shares, or comments—represents not just a digital footprint but also the genuine affection of a real person behind the screen. As CF37 noted, ”We, all the fans, including core fans like myself, partake in this data labor purely out of love, without any financial or other incentives. Our sole aim is to elevate our idol’s stature.”
@@
4.3.1. Convincing Skeptical General Fans
@@
As GF4, a general fan with two years of experience whom we interviewed, put it,
”Participating in algorithmic actions is really exhausting and complicated. I have to figure out what I need to do on each platform, and I spend a lot of time browsing posts and even playing videos, which drains my energy and ties up my devices. Sometimes, I don’t even know what my actions mean—does my small effort as a single user really make a difference? And if I don’t do it, there will be other fans who will.” (GF4)
@@
For instance, core fans will argue that if their beloved idol’s data performance is outstanding, similar benefits achieved by predecessors can be expected. If the benefits are not immediately evident from past cases, core fans will highlight the authoritative nature of the platform where they engage in data labor. They also emphasize the prominence of idols who are part of the same data performance competition, thereby crafting a narrative that suggests brands and TV programmers closely monitor this data. This way, core fans imply to the general fans that their idol could gain significant resource boosts through this data and their collective labor efforts.
4.3.2. Making Algorithmic Manipulation Tutorials as Simple as Possible

### v2

@@ 5. Findings: Fan-Driven Folk Theories: The Foundation for Guiding Collective Algorithmic Actions (RQ2)
5.1. Fan-Driven Methods in Understanding Algorithms
Before launching collective actions targeting algorithms, fans must have a method to thoroughly understand the algorithms, allowing them to perceive whether their actions can effectively influence them and adjust their strategies accordingly based on this assessment. Core fans use the term ”decode” to describe the initial step in developing their understanding of algorithms. They see the algorithm as a black box; users are unaware of its exact operational logic and can only decode it through continuous trial and error. Eventually, this user-driven process creates a set of tutorials that can genuinely influence the algorithm’s output in fans’ perception.
@@ 5.1. Fan-Driven Methods in Understanding Algorithms
To develop a clear understanding of the algorithm, initially, core fans often gather basic tutorials from other fan groups, as these groups also engage in decoding algorithms and instructing their fans. This is because all fan groups are actively decoding various social media algorithms to increase the visibility of their idols. To engage more general fans, these algorithm manipulation tutorials are always made public. As previously mentioned, there is already a consensus among core and general fans that various interactions on social media impact their idols. Then, it is important to understand which specific interactions can most effectively leverage recommendation algorithms to maximize the visibility of their idols.
@@
By observing patterns in algorithm behavior from various idol fanbases, core fans can formulate an initial strategy for their own group. For example, CF26 and her core fan colleagues examine tutorials from different idol fanbases, selecting and integrating them based on their experience. The core fan then guides general fans in improving the idol’s visibility through continual experimentation, eventually tailoring an algorithmic tutorial that best fits their fanbase. This is because fan groups are heterogeneous. The behavior and characteristics of fans from different idol groups can vary hugely, so core fans must carefully understand and adapt to the style of their corresponding general fan group. Moreover, since these are folk understandings, the algorithmic tutorials created by other fan groups are not necessarily effective. Core fans are often skeptical of tutorials from other groups and prefer to develop their own understanding through repeated experimentation and testing.
@@
Improvements in data performance are often perceptible only to core fans, rather than being quantifiable through objective social media metrics. They observe whether the manipulations lead to higher comment rankings and visibility than other fanbases. Recently, CF26 got information from a fan working at Sina (Weibo company) about how to make a topic become a trending topic. When a topic related to an idol enters the trending section by the recommended algorithm, it gains visibility, making this achievement extremely important for fans.
@@
Within a single topic, there are often many posts, but how to intervene so that the algorithm recommends the topic and pushes it into Weibo’s trending list, thereby gaining greater visibility, is a common concern for fans. CF26 mentioned that the new information from the Sina staff asks the fans to engage with the post displayed on the top of the topic, called “source of trend.” This means that interacting with the top posts within a topic, by giving them more retweets, likes, and comments, will make the topic more likely to enter the trending section rather than interacting with other, less prominent posts. When CF26 and her followers tried to make a new trend topic for their idol, they utilized this information to make a new tutorial, and engaged with the “source of trending.” After the topic successfully becomes a new trending topic, CF26 uploads a new post: “The new tutorial is so useful! It has been trending for about two hours!” CF26 told us that she uses this period as a verification period to ensure the new tutorial makes sense.
@@
5.2. Expand Folk Understanding of Algorithms across Platforms in Different Nations and across Different Fan Bases
If such collective action involves multiple countries, fans must be able to understand the differences across these countries. The essence of these fan algorithmic practices lies in a mindset centered on “algorithmic decoding” according to our participants, which guides fans’ collective actions across various platforms and countries. This means that, regardless of location, fans are essentially trying to understand the operational logic of the algorithm in order to manipulate it in reverse, increasing the visibility of content related to their idol. This approach is not confined to a single platform or nation; fans worldwide, particularly K-pop fans, apply this mindset to collectively influence data related to their idols on almost all platforms, often generating significant impact. For instance, on social media platforms like X (formerly Twitter), Instagram, and Weibo, fans might launch large-scale campaigns where they post the same hashtag en masse to boost its visibility or extensively like their idols’ posts to demonstrate popularity. On video platforms with recommendation algorithms, such as YouTube and BiliBili, fans organize targeted efforts to like their idols’ videos and repeatedly play them in specific ways to inflate view counts. Similarly, on short video platforms like TikTok, fans engage in activities such as liking, commenting, and fully watching their idols’ videos to enhance their recommendation metrics. Even on niche voting apps, fans often participate in repetitive and highly organized voting campaigns to support their idols. This fan- driven algorithmic practice is not restricted by platform or national boundaries. Instead, it reflects an internalized mindset within fan communities, driving organized collective actions that are applied broadly across the internet to amplify their idols’ presence and influence.
@@
If this collective action involves different groups of fans, it is essential to understand the unique characteristics of these groups and leverage these traits to effectively organize them. Consequently, the interpretation of algorithms and the formulation of tutorials for algorithm manipulation by core fans vary between cultures and fanbases, and are largely based on empirical observations.
In the context of the entertainment industry and social media, Chinese fans often employ a specific ”nested comment (buildings within buildings)” strategy, which means comments within comments, a hierarchical commenting method used to manipulate data performance. CF26 highlights the unique dynamics between Chinese K-pop fans and fans of Chinese celebrities across social media platforms, noting: ”The habits and scale of Chinese K-pop fans differ from those of fans of Chinese celebrities. They are smaller in scale and easier to manage.” In China, the ”die-hard fan” status is a feature on Weibo that indicates a high level of interaction between a fan and a specific content bloggers. To obtain this status, a fan must interact with the bloggers for at least five days within the past 30 days. This status must be acquired separately for each blogger. The ”nested comment” strategy in China is influenced by users’ ”die-hard fan” status and Weibo VIP membership. CF26 explains that if users meet both conditions, their fan community’s comments will gain higher visibility. However, if they lack ”die-hard fan” status, this strategy may actually have a negative impact.
Furthermore, CF26 details that most general fans do not meet the criteria for either ”die-hard fan” status or VIP membership. The ”die-hard fan” status represents high-frequency interactions with a specific blogger, while obtaining VIP membership requires monetary contributions. General fans also find it difficult to maintain ”die-hard fan” status across multiple bloggers simultaneously. As a result, the ”nested comment” strategy could reduce visibility if it is employed by users who fail to meet these two requirements.
In this context, the term ”blogger” here broadly refers to all Weibo users who post content, not just professional content creators or influencers. Essentially, any individual who publishes content on Weibo can be considered a ”blogger.” For general fans, maintaining ”die-hard fan” status across multiple bloggers is challenging, as it requires consistent effort and resources. Consequently, the ”nested comment” strategy intended to amplify visibility on social media may backfire if participants do not meet the ”die-hard fan” and VIP requirements. This could lead to a decrease in the visibility of collective algorithmic actions, as these strategies heavily rely on the contributions of highly engaged and financially supportive fans.
CF43 agrees that for fans of Chinese celebrities, core fans often advise general fans against using the ”nested comment” strategy. CF43 elaborates: ”This is because Chinese celebrities have a large and uncontrollable fan base on Chinese social media. Directly instructing all fans to avoid using nested comments is a simpler and more efficient approach. When general fans, especially those who do not meet the aforementioned criteria, attempt to use this strategy, the platform may classify their actions as ’bot-like’ behavior, considering it a disruption to social media order. Once flagged as ’bot activity,’ fans’ comments will be downgraded and pushed to the bottom of the comment section, losing all visibility and rendering previous efforts futile.”
In contrast, Chinese K-pop fans, due to their smaller scale and more clustered activities, enable core fans to exert greater control over general fans, permitting only those with ”die-hard fan” status and VIP membership to participate in ”nested comments.” The effectiveness of this strategy among Chinese K-pop fans highlights the different ways fan communities adapt to the same platform algorithms. Despite facing similar requirements for ”die-hard fan” and VIP membership, fans of Chinese celebrities, being more difficult to manage, opt to ban all fans from using ”nested comments,” whereas K-pop fans can better coordinate their fanbase, allowing them to leverage the ”nested comment” strategy.
@@
CF11, who manages a data group in an idol’s fan club, also highlighted how fans from different countries, even those of the same idol, employ distinct strategies on platforms like YouTube. CF11 noted that view counts of fancams on YouTube are crucial for K-pop idols, and K-pop fans worldwide, regardless of their country, strive to increase these counts. However, CF11 pointed out a stark contrast in techniques: while international fan organizations advise fans to remain logged in while viewing, Chinese fans do not recommend this approach. CF11 further explained their understanding of YouTube’s view count algorithm: “YouTube’s algorithm limits the number of view counts to 3-5 plays per video per day from a single user. Users logged in can’t contribute more than five views. Therefore, to exceed this limit, one should not log in.” CF11 believes these different approaches stem from the varying user habits of the two fan bases, “Chinese fans are used to repeatedly engaging in data labor, often exceeding five views per day, hence our preference for not logging in. In contrast, the international fan base, predominantly from Europe and the U.S., tends to view 3-5 times daily. For them, ensuring each view is counted becomes vital, making logged-in viewing more appropriate.”
These differences between fan communities also suggest that if core fans do not develop a precise understanding of platform algorithms, their collective actions are unlikely to succeed. Without such knowledge, they cannot determine which behaviors will be effective or ineffective, and may even inadvertently trigger negative consequences.
@@
5.3. Expand Folk Understanding of Algorithm across Different Levels of Algorithm Transparency
@@
Since this is a collective action targeting algorithms, it must take into account the differences between the algorithms of various platforms, particularly in terms of transparency. Effectively leveraging platform transparency can enhance the efficiency of the action. Even when dealing with platforms with low transparency, fans must strive to understand the underlying algorithmic mechanisms. While deciphering algorithms can be intricate and challenging, as seen with platforms like Weibo, some platforms offer more transparency. CF39 mentioned that many platforms provide official explanations of their recommendation systems, simplifying the process for core fans to create tutorials.
@@
For example, BiliBili, a popular platform in China, officially states that the same video can be counted 3-5 times per day per person for play data. Although full playback isn’t necessary for counting views, the completion rate in playing is a perceived important factor in the recommendation algorithm. This clarity gives fans straightforward tutorials for contributing to BiliBili’s play count.
@@
However, according to our participants, the criteria for BiliBili’s trending list are less clear for fans. Within the Chinese fandom, it was rumored that a video must achieve 100k plays within 24 hours, along with a specific number of likes, shares, comments, and other metrics, to feature on the trending list algorithmically. Yet, there are instances where videos have made it to the trending list with fewer than 100k plays or exceeded this threshold without being featured. CF39 expressed frustration over this ambiguity, stating, ”We understand that activities like plays, likes, favorites, comments, and shares positively impact trending potential, but the exact mechanism is unclear. We cannot definitively influence the recommendation algorithm. Our only recourse is for fans to engage in data labor blindly and keep experimenting.” (CF39)
@@
We have described how core fans gradually develop a folk understanding of how algorithms operate on different platforms. It is important to note that while these understandings are not entirely conveyed to general fans, they serve as a crucial foundation for core fans to develop algorithm manipulation tutorials. All the core fans we interviewed acknowledged that without a clear understanding of platform algorithms, collective algorithmic action would be impossible. CF34, a core fan with over ten years of experience, explained to us, ”Ten years ago, we lacked understanding of algorithms and the idea of collective algorithmic action. Fans didn’t unite. Even after recognizing the importance of recommendation algorithms, it took time to learn how to coordinate effectively. Without clear direction, mobilizing millions of fans would have been impossible. Now, our actions are organized and disciplined, and fans know precisely what is expected of them, resulting in a observed impact.”
@@
6. Findings: “Enable the Foolish to Learn, Motivate the Lazy to Act”: The Principles of Mobilizing Millions of Fans (RQ3)
@@
CF37 challenged the previous Chinese media’s portrayal of fans as a highly efficient, organized group that immediately responds to requests for collective algorithmic action. According to CF37, this depiction is misleading. CF37 explained that each act of data engagement, likes, shares, or comments, represents not just a digital footprint but also the genuine affection of a real person behind the screen. As CF37 noted, ”We, all the fans, including core fans like myself, partake in this data labor purely out of love, without any financial or other incentives. Our sole aim is to elevate our idol’s stature.”
@@
6.1. Convincing Skeptical General Fans
@@
As GF4, a general fan with two years of experience whom we interviewed, put it, ”Participating in algorithmic actions is really exhausting and complicated. I have to figure out what I need to do on each platform, and I spend a lot of time browsing posts and even playing videos, which drains my energy and ties up my devices. Sometimes, I don’t even know what my actions mean. Does my small effort as a single user really make a difference? And if I don’t do it, there will be other fans who will.”
@@
For instance, core fans will argue that if their beloved idol’s data performance is outstanding, similar benefits achieved by predecessors can be expected. If the benefits are not immediately evident from past cases, core fans will highlight the authoritative nature of the platform where they engage in data labor. They also emphasize the prominence of idols who are part of the same data performance competition, thereby crafting a narrative that suggests brands and TV programmers closely monitor this data. This way, core fans imply to the general fans that their idol could gain resource boosts through this data and their collective labor efforts.
6.2. Making Algorithmic Manipulation Tutorials as Simple as Possible